OpenStudy (anonymous):

solve the equation for z .12z+5=15z+14

OpenStudy (anonymous):

@Parvathysubhash @eccme @Diyadiya

OpenStudy (anonymous):

12z-15z=14-5 -3z=9 z=9/-3 z=-3

OpenStudy (anonymous):

15z-12z=5-14,3z=-9,z=-3

OpenStudy (anonymous):

yhanksss guyss

OpenStudy (anonymous):

ur welcum!!

OpenStudy (anonymous):

wc if u dnt mind u giv me and eccme a medal

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

how do u do that?

OpenStudy (anonymous):

whtt

OpenStudy (anonymous):

k next question

OpenStudy (anonymous):

k

OpenStudy (anonymous):

dats mean ! u took away mine!!:( just goofin,dont mind it!!